# When Federated Learning Meets Quantum Computing: Survey and Research Opportunities

### v2

@@ 6 FL on NISQ Devices
$\bullet$ Challenges: While offering quantum computational advantages, NISQ devices are inherently constrained by high levels of noise [80, 81], limited qubit count [82, 83], and low gate fidelity [84, 85]. The primary challenge in utilizing these devices within an FL framework is managing quantum noise. Quantum states, especially in multi-qubit systems, are highly sensitive to decoherence, leading to loss of information and reduced model accuracy [23]. Additionally, the relatively small number of qubits limits the complexity of quantum computations, restricting the depth of quantum circuits involved in federated tasks. In the near term, error mitigation techniques rather than full quantum error correction may be the most viable route forward to realize FL on NISQ devices [86].
@@
[87] Demonstrates adaptive methods that improve convergence in nonconvex FL Sensitive to heterogeneity, client drift not fully resolved Moderate [adaptive optimization] Synchronous aggregation
[88] Integrates Pareto optimization ensuring fairness and robustness Complexity in multi-objective trade-offs can slow convergence Moderate [multi-objective balance] Asynchronous aggregation
[52] Faster convergence and reduced communication rounds in QFL Limited generalization to complex datasets with significant noise High [variational circuit] Synchronous aggregation
[64] Comprehensive review of QAOA and its variants; hardware-focused insights No practical implementation insights for distributed systems High [optimized circuit depth] Hybrid aggregation
[61] Effective for time-series data processing with reduced training time Limited robustness in non-time-series contexts High [efficient QLSTM structure] Asynchronous aggregation
[77] Proposes hybrid quantum-classical training; preserves data privacy Scalability issues for larger quantum data sets Moderate [hybrid circuit integration] Synchronous aggregation
[89] Integrates QCNNs for non-IID data in medical imaging; reduces communication rounds Scalability and noise resilience not extensively addressed High [shallow quantum circuits] Synchronous aggregation
[19] Adaptable across quantum hardware; strong results on multiple datasets Accuracy fluctuations due to noise and gate fidelity, limited large-scale insights Moderate [adaptive qubit usage] Asynchronous aggregation
@@ 6.1 Optimization Techniques for QFL
To overcome the inherent limitations of NISQ devices, various optimization strategies have emerged. Central to this effort are VQAs, which operate by embedding quantum circuits within a classical optimization loop. In this framework, quantum circuits explore a vast solution space, while classical processors adjust variational parameters based on performance evaluations. The classical optimizer iteratively refines the parameters until a convergence criterion is met, ensuring that the workload is effectively shared between the quantum and classical components [87]. Such a hybrid approach has been crucial for leveraging quantum advantages without overwhelming the capacity of NISQ devices. However, optimization in hybrid systems can be resource-intensive. Using ML, the authors in [90] attempted to predict optimal quantum parameters with QAOA. Remarkably, their technique reduces the number of iterations in the optimization loop by as much as 65.7%. Instead of initiating each optimization loop from scratch, they harnessed correlations between lower-depth and higher-depth QAOA parameters. This intelligent initialization significantly reduces computational cost, accelerating convergence. Although not yet applied to FL, this system appears promising in QFL.
@@
Beyond classical ML, there is a growing focus on enhancing distributed quantum systems. QFL tackles optimization challenges in decentralized and heterogeneous environments. Techniques like FedMGDA+ [88] are designed to ensure fairness and robustness, even when faced with adversarial conditions and uneven data distribution. The adaptability of these methods makes them particularly valuable for ensuring stability in real-world quantum networks. In addition, optimization benefits from TL strategies, which accelerate the training process by transferring knowledge from previously optimized quantum systems. QFL researchers have shown that the transfer of parameters from simpler quantum tasks can speed up optimization for complex problems, reducing the overall computational burden [77]. This cross-task learning is pivotal in accelerating multitask optimization within quantum systems. In essence, these quantum-classical optimization techniques spanning from VQAs to FL and TL are steadily transforming the landscape of QC.
@@
An evaluation of $\mathrm{QUE}$ for optimization‐focused QFL methods shows that VQA‐based schemes like the adaptive optimizer in [87] typically employ moderate‐depth circuits (e.g. $\mathrm{Q}_{alloc}=5$, $L=4$) yielding $\mathrm{QUE}\approx 1.0$, which aligns with their preference for synchronous global updates to exploit full‐batch curvature corrections. In contrast, Pareto‐based multi‐objective optimizers such as FedMGDA+ [88] often run on slightly shallower ansatze ($\mathrm{Q}_{alloc}=4$, $L=3$), giving $\mathrm{QUE}\approx 0.80$ and benefiting from asynchronous local updates that accommodate imbalanced data distributions. TL accelerators like the parameter‐reuse strategy in [77] sit between these extremes with $\mathrm{QUE}\approx 0.90$ and favor hybrid classical–quantum fusion to blend pretrained classical weights with fresh quantum gradients. Beyond the optimization benefits, these methods must address the overhead of metric‐tensor estimation in  [52], the staleness introduced by lock‐free updates in [88], and the calibration drift that arises when classical weights are merged with quantum‐tuned parameters in transfer scenarios.
@@ 6.2 Hybrid Gradient Descent Techniques for QFL
Gradient descent is at the heart of most ML algorithms, and it plays a pivotal role in updating parameters across distributed nodes in FL. As quantum devices become more integrated with FL systems, hybrid gradient descent techniques have emerged. These techniques leverage the computational power of quantum processors to compute partial gradients or assist in optimizing specific model parameters. The combination of quantum and classical resources accelerates convergence by efficiently partitioning the tasks between quantum and classical nodes [91].
@@
A notable research in [52] introduced Federated Quantum Natural Gradient Descent (FQNGD), a novel optimization method for QFL, utilizing VQCs to enhance convergence and reduce communication overhead. The experimental results on the MNIST dataset showed superior accuracy (more than 99% for binary classification) over traditional SGD methods. This technique markedly enhances convergence rates, especially within the realm of NISQ devices [92]. Furthermore, by allowing quantum devices to compute gradients for only a subset of parameters, hybrid systems can dynamically balance computational loads and reduce training bottleneck [93]. However, reliance on block-diagonal approximations of the Fubini-Study could limit robustness, as it may overlook complex entanglements in deeper circuits, impacting practical performance.
@@
Analysis of $\mathrm{QUE}$ for hybrid gradient methods shows that FQNGD’s variational circuits ($\mathrm{Q}_{alloc}=5$, $L=4$) achieve $\mathrm{QUE}\approx 1.0$, matching its preference for synchronous aggregation to exploit full‐batch curvature corrections via the quantum Fisher metric [52]. Asynchronous update schemes, exemplified by the Paillier‐secured protocol in [53], typically run on shallower ansatz ($\mathrm{Q}_{alloc}=4$, $L=3$), yielding $\mathrm{QUE}\approx 0.80$ and benefiting from low-latency, lock‐free gradient pushes. Fusion strategies that combine classical FedAvg with quantum natural gradients, such as the hybrid model in [54] operate at moderate QUE ($75\%$$85\%$) and balance robustness and acceleration by tuning a mixing coefficient. Excluding common measurement and communication overheads, these methods face the challenge of efficiently estimating and inverting high-dimensional metric tensors on noisy hardware, and of dynamically adjusting aggregation weights to prevent one modality from dominating the learning trajectory.
@@ 6.3 Quantum-Inspired Classical Algorithms
With the introduction of quantum neural computation by [94], QNN received significant attention from QFL community. This pioneering theoretical framework is a crucial advancement in the integration of quantum computation concepts with traditional neural network models, setting the stage for subsequent studies of quantum-inspired classical algorithms. Building on this foundation, numerous quantum-enhanced classical techniques have been developed, each targeting particular computational constraints of classical systems while utilizing quantum benefits.
Building on the challenges of quantum circuit stability, FedQNN [19] presents a more generalized federated quantum framework for various datasets, including genomics and healthcare. A critical strength of the framework is its adaptability to different quantum hardware architectures, evidenced by evaluations on IBM QPUs. The authors in [89] introduced a novel FL framework using QCNNs for privacy-preserving collaborative learning, particularly suited for non-IID medical data distributions. The novelty lies in the use of quantum circuits to enhance feature extraction and classification within a federated setting. The development of the Quantum Long Short-Term Memory (QLSTM) [61] network represents a significant advancement aimed at overcoming the challenges faced by classical LSTMs [95], especially in managing long-term dependencies in time-series data. The unique design of the QLSTM employs quantum operations to bolster memory retention while minimizing the computational demands typical of classical LSTMs. Further, Quantum Fuzzy Neural Networks (QFuzzyNNs) [60], inspired by tensor network methodologies, present a unique combination of quantum computing principles and fuzzy logic. By utilizing quantum superposition and entanglement, QFuzzyNNs conduct fuzzy inference, enhancing decision-making capabilities even with incomplete information [96]. In Fig. 8, we present the advantages and disadvantages of various neural networks when combined with quantum capabilities.
@@
On a slightly different track, QAOA [97] has simultaneously established itself as a potent method to address combinatorial optimization problems. QAOA functions by alternating between quantum and classical optimization stages, enabling it to effectively search through vast solution spaces typically beyond the reach of classical algorithms. The principal innovation of QAOA is its application of quantum circuits for state evolution, followed by classical updates to refine the variational parameters [64]. This hybrid method has demonstrated remarkable potential in solving the MaxCut problem, consistently surpassing classical algorithms in both convergence speed and solution quality. The capacity of QAOA to provide near-optimal solutions for NP-hard problems underscores its importance in diverse optimization tasks, ranging from logistics [98, 99] to cryptography [100, 101]. Recently, researchers have explored the application of quantum tensors for cyber-physical-social systems [62, 63] and revealed that tensor decomposition techniques can substantially decrease communication overhead in distributed learning settings.
An inspection of $\mathrm{QUE}$ for quantum-inspired classical algorithms reveals varied occupancy levels tied to circuit depth and qubit usage. FedQNN’s generalized framework ($\mathrm{Q}_{alloc}=5$, $L=3$) attains $\mathrm{QUE}\approx 1.0$, making it well suited to asynchronous updates that exploit its high utilization and low-latency gradient pushes [19]. The QCNN-based privacy model ($\mathrm{Q}_{alloc}=4$, $L=2$) yields $\mathrm{QUE}=1.0$ and favors synchronous aggregation to ensure consistent feature extraction across non-IID medical nodes [89]. QLSTM’s deeper memory-enhancing circuits ($\mathrm{Q}_{alloc}=4$, $L=4$) report $\mathrm{QUE}\approx 1.0$ and employ specialized asynchronous updates to balance sequence processing with qubit coherence constraints [61]. Despite these strengths, tensor-network hybrids ($\mathrm{Q}_{alloc}=4$, $L=5$) show slightly reduced QUE ($\approx 0.80$) and require hybrid classical–quantum fusion to reconcile their expressive power with aggregation consistency [63]. Limits unique to these hybrid frameworks include the overhead of repeated state re-preparation for classical data encoding and the absence of standardized benchmarks for comparing QUE across heterogeneous quantum-inspired pipelines.

### v4

@@ 6 FL on NISQ Devices
$\bullet$ Challenges: While offering quantum computational advantages, NISQ devices are inherently constrained by high levels of noise [cao2021nisq, yetis2021investigation], limited qubit count [tannu2019not, nash2020quantum], and low gate fidelity [chen2023high, berrios2012high]. The primary challenge in utilizing these devices within an FL framework is managing quantum noise. Quantum states, especially in multi-qubit systems, are highly sensitive to decoherence, leading to loss of information and reduced model accuracy [Preskill2018]. Additionally, the relatively small number of qubits limits the complexity of quantum computations, restricting the depth of quantum circuits involved in federated tasks.
Most quantum systems use additional qubits for Quantum Error Correction (QEC) techniques. Some early-stage techniques, such as repetition codes and surface codes, provide partial mitigation of noise in NISQ systems [khalid2024quantumnetwork]. However, their scaling characteristics pose distinct challenges in FL. Each client node operating under QEC must allocate additional qubits for encoding logical qubits, increasing $Q_{\text{alloc}}$ well beyond the original model requirements. In large-scale FL deployments, this qubit overhead compounds across clients, limiting the feasible number of participating devices. In asynchronous settings, heterogeneous QEC performance across clients can lead to model inconsistency if updates are based on variably protected computations. These constraints imply that while early QEC can enhance noise resilience, its direct adoption in FL must be balanced against the cost in qubit resources, communication time, and aggregation stability. It indicates that error mitigation techniques, rather than full QEC, may be the most viable route forward to realize FL on NISQ devices [ErrorMitigation].
Furthermore in practical QFL deployments, clients rarely have identical logical-qubit capacities, coherence times, or noise characteristics, and a rigid assumption of a single uniform PQC architecture would either exclude weaker devices or inject unstable gradients that slow or destabilize convergence [QFL-TL-2]. A more realistic strategy treats hardware variation as a structured design constraint i.e., the server specifies a nested family of ansatzes in which low-depth, low-qubit subcircuits constitute a shared core and deeper blocks extend this core for clients with larger logical budgets, analogous in spirit to model-heterogeneous FL, but adapted to parametrized quantum circuits [diao_heterofl_2021]. Each client then instantiates the deepest sub-ansatz compatible with its logical-qubit and shot budget, updates only the parameters it supports, and contributes these to aggregation, while the coordinator averages each parameter block over the subset of clients that implements it [diao_heterofl_2021, han_quorus_2025]. This perspective aligns with recent QFL frameworks that employ layerwise or blockwise objectives to handle varying circuit depths, where heterogeneous-depth clients optimize different truncations of a shared PQC and still produce a coherent global model [han_quorus_2025]. Extremely constrained clients can still play meaningful roles by hosting shallow encoders, measurement heads, or hybrid split-learning frontends that interface with quantum backends. This verifies that limited local quantum footprint can be combined with stronger remote resources while preserving data locality [cowlessur_hqsl_2025]. Combined with the temporal mechanisms, in which asynchronous or semi-synchronous aggregation assigns staleness and depth-aware weights to updates, these constructions ensure that devices with diverse logical-qubit resources contribute systematically rather than sporadically
@@
[QFL-Optimization] Demonstrates adaptive methods that improve convergence in nonconvex FL Sensitive to heterogeneity, client drift not fully resolved Moderate [adaptive optimization] Synchronous aggregation
[QFL-MultiOptimization] Integrates Pareto optimization ensuring fairness and robustness Complexity in multi-objective trade-offs can slow convergence Moderate [multi-objective balance] Asynchronous aggregation
[QFL-QNGD] Faster convergence and reduced communication rounds in QFL Limited generalization to complex datasets with significant noise High [variational circuit] Synchronous aggregation
[QFL-QAOA] Comprehensive review of QAOA and its variants; hardware-focused insights No practical implementation insights for distributed systems High [optimized circuit depth] Hybrid aggregation
[QFL-QLSTM] Effective for time-series data processing with reduced training time Limited robustness in non-time-series contexts High [efficient QLSTM structure] Asynchronous aggregation
[QFL-TL-2] Proposes hybrid quantum-classical training; preserves data privacy Scalability issues for larger quantum data sets Moderate [hybrid circuit integration] Synchronous aggregation
[QFL-QCNN] Integrates QCNNs for non-IID data in medical imaging; reduces communication rounds Scalability and noise resilience not extensively addressed High [shallow quantum circuits] Synchronous aggregation
[QFL2] Adaptable across quantum hardware; strong results on multiple datasets Accuracy fluctuations due to noise and gate fidelity, limited large-scale insights Moderate [adaptive qubit usage] Asynchronous aggregation
@@ 6.1 Optimization Techniques for QFL
To overcome the inherent limitations of NISQ devices, various optimization strategies have emerged. Central to this effort are VQAs, which operate by embedding quantum circuits within a classical optimization loop. In this framework, quantum circuits explore a vast solution space, while classical processors adjust variational parameters based on performance evaluations. The classical optimizer iteratively refines the parameters until a convergence criterion is met, ensuring that the workload is effectively shared between the quantum and classical components [QFL-Optimization]. Such a hybrid approach has been crucial for leveraging quantum advantages without overwhelming the capacity of NISQ devices. However, optimization in hybrid systems can be resource-intensive. Using ML, the authors in [ML-Optimization] attempted to predict optimal quantum parameters with QAOA. Remarkably, their technique reduces the number of iterations in the optimization loop by as much as 65.7%. Instead of initiating each optimization loop from scratch, they harnessed correlations between lower-depth and higher-depth QAOA parameters. This intelligent initialization significantly reduces computational cost, accelerating convergence.
@@
Beyond classical ML, there is a growing focus on enhancing distributed quantum systems. QFL tackles optimization challenges in decentralized and heterogeneous environments. Techniques like FedMGDA+ [QFL-MultiOptimization] are designed to ensure fairness and robustness, even when faced with adversarial conditions and uneven data distribution. The adaptability of these methods makes them particularly valuable for ensuring stability in real-world quantum networks. QFL researchers have shown that the transfer of parameters from simpler quantum tasks can speed up optimization for complex problems, reducing the overall computational burden [QFL-TL-2]. This cross-task learning is pivotal in accelerating multitask optimization within quantum systems.
@@
An evaluation of $\mathrm{QUE}$ for optimization‐focused QFL methods shows that VQA‐based schemes like the adaptive optimizer in [QFL-Optimization] typically employ moderate‐depth circuits (e.g. $\mathrm{Q}_{alloc}=5$, $L=4$) yielding $\mathrm{QUE}\approx 1.0$, which aligns with their preference for synchronous global updates to exploit full‐batch curvature corrections. In contrast, Pareto‐based multi‐objective optimizers such as FedMGDA+ [QFL-MultiOptimization] often run on slightly shallower ansatze ($\mathrm{Q}_{alloc}=4$, $L=3$), giving $\mathrm{QUE}\approx 0.80$ and benefiting from asynchronous local updates that accommodate imbalanced data distributions. TL accelerators like the parameter‐reuse strategy in [QFL-TL-2] sit between these extremes with $\mathrm{QUE}\approx 0.90$ and favor hybrid classical–quantum fusion to blend pretrained classical weights with fresh quantum gradients. Beyond the optimization benefits, these methods must address the overhead of metric‐tensor estimation in  [QFL-QNGD], the staleness introduced by lock‐free updates in [QFL-MultiOptimization], and the calibration drift that arises when classical weights are merged with quantum‐tuned parameters in transfer scenarios.
@@ 6.2 Hybrid Gradient Descent Techniques for QFL
Gradient descent is at the heart of most ML algorithms, and it plays a pivotal role in updating parameters across distributed nodes in FL. As quantum devices become more integrated with FL systems, hybrid gradient descent techniques have emerged. These techniques leverage the computational power of quantum processors to compute partial gradients or assist in optimizing specific model parameters. The combination of quantum and classical resources accelerates convergence by efficiently partitioning the tasks between quantum and classical nodes [optimization2].
@@
A notable research in [QFL-QNGD] introduced Federated Quantum Natural Gradient Descent (FQNGD), a novel optimization method for QFL, utilizing VQCs to enhance convergence and reduce communication overhead. The experimental results on the MNIST dataset showed superior accuracy (more than 99% for binary classification) over traditional SGD methods. This technique markedly enhances convergence rates, especially within the realm of NISQ devices [optimization3]. Furthermore, by allowing quantum devices to compute gradients for only a subset of parameters, hybrid systems can dynamically balance computational loads and reduce training bottleneck [optimization4]. However, reliance on block-diagonal approximations of the Fubini-Study could limit robustness, as it may overlook complex entanglements in deeper circuits, impacting practical performance.
@@
Analysis of $\mathrm{QUE}$ for hybrid gradient methods shows that FQNGD’s variational circuits ($\mathrm{Q}_{alloc}=5$, $L=4$) achieve $\mathrm{QUE}\approx 1.0$, matching its preference for synchronous aggregation to exploit full‐batch curvature corrections via the quantum Fisher metric [QFL-QNGD]. Asynchronous update schemes, exemplified by the Paillier‐secured protocol in [QFL-AsynchAgg], typically run on shallower ansatz ($\mathrm{Q}_{alloc}=4$, $L=3$), yielding $\mathrm{QUE}\approx 0.80$ and benefiting from low-latency, lock‐free gradient pushes. Fusion strategies that combine classical FedAvg with quantum natural gradients, such as the hybrid model in [FQE2] operate at moderate QUE ($75\%$$85\%$) and balance robustness and acceleration by tuning a mixing coefficient. However, these methods still face the challenge of efficiently estimating and inverting high-dimensional metric tensors on noisy hardware.
$\bullet$ Hybrid gradient implementation: To optimally implement hybrid gradient descent in a QFL setting, it is important to design it such that it respects both the characteristics of PQCs and the practical constraints of distributed FL training. Let us partition the global parameter vector into quantum and classical components as $\boldsymbol{\theta}\;=\;(\boldsymbol{\theta}_{Q},\,\boldsymbol{\theta}_{C}),$ reflecting the fused architecture [Mitarai2018QCL, Schuld2019AnalyticGradients]. This is because PQC parameters are updated through measurement-based estimators that differ fundamentally from the backpropagation-based updates applied to classical layers. Each client $k$ then evaluates its local objective $\mathcal{L}_{k}$ and constructs a hybrid gradient that blends quantum and classical components:
$\nabla\mathcal{L}_{k}(\boldsymbol{\theta})\;=\;\alpha\,\nabla_{\boldsymbol{\theta}_{Q}}\mathcal{L}_{k}(\boldsymbol{\theta})\;+\;(1-\alpha)\,\nabla_{\boldsymbol{\theta}_{C}}\mathcal{L}_{k}(\boldsymbol{\theta}),$
where the coefficient $\alpha\in[0,1]$ regulates the relative contribution of quantum updates. This convex weighting provides flexibility, allowing the update dynamics to be tuned in response to hardware noise levels, QPU queue times, or task-specific sensitivity to quantum features. For the quantum component, gradients are estimated via the parameter-shift rule, which for a parameter $\theta_{Q,i}$, evaluated at 2 shifted points $\pm s$ takes the form
$\frac{\partial\mathcal{L}_{k}}{\partial\theta_{Q,i}}=\frac{1}{2}\!\left[\mathcal{L}_{k}(\boldsymbol{\theta}_{Q}+s\,\mathbf{e}_{i},\,\boldsymbol{\theta}_{C})-\mathcal{L}_{k}(\boldsymbol{\theta}_{Q}-s\,\mathbf{e}_{i},\,\boldsymbol{\theta}_{C})\right]$
with shift $s=\pi/2$ and $\mathbf{e}_{i}$ the $i$th standard basis vector [Schuld2019AnalyticGradients, Crooks2019ParamShift]. This rule is compatible with a wide range of NISQ devices, avoids the overhead of full adjoint differentiation, and introduces minimal circuit depth change, thereby preserving coherence times. As discussed above, when applicable, the PQC gradient can be preconditioned using the quantum natural gradient. This involves estimating the Fubini–Study metric $\mathbf{F}_{k}$ and rescaling the gradient via $\widetilde{\nabla}_{\boldsymbol{\theta}_{Q}}\mathcal{L}_{k}\;=\;\mathbf{F}_{k}^{-1}\,\nabla_{\boldsymbol{\theta}_{Q}}\mathcal{L}_{k},$ which aligns the update direction with the local geometry of the Hilbert space. This approach mitigates the effects of barren plateaus and poor conditioning, enabling faster and more stable convergence. The local update is then performed with learning rate $\eta$ as $\boldsymbol{\theta}^{\,t+1}_{k}=\boldsymbol{\theta}^{\,t}-\eta\,\nabla\mathcal{L}_{k}(\boldsymbol{\theta}^{\,t}),$ or using the preconditioned gradient in the quantum block if natural gradient correction is applied. Conducting multiple local steps before communication balances QPU access and reduces communication overhead in federated setups. Aggregation at the server can proceed synchronously, where sample-proportional weights $p_{k}=n_{k}/\sum_{j\in\mathcal{S}_{t}}n_{j}$ yield $\boldsymbol{\theta}^{\,t+1}=\boldsymbol{\theta}^{\,t}-\eta\sum_{k\in\mathcal{S}_{t}}p_{k}\,\nabla\mathcal{L}_{k}(\boldsymbol{\theta}^{\,t}),$ ensuring that quantum and classical parameters evolve together across clients. Alternatively, in asynchronous settings, stale updates can be down-weighted via a decay factor $w_{k}(\Delta t_{k})\in(0,1]$ dependent on staleness $\Delta t_{k}$:
$\boldsymbol{\theta}^{\,t+1}=\boldsymbol{\theta}^{\,t}-\eta\sum_{k\in\mathcal{A}_{t}}p_{k}\,w_{k}(\Delta t_{k})\,\Big[\alpha\,\widetilde{\nabla}_{\boldsymbol{\theta}_{Q}}\mathcal{L}_{k}+(1-\alpha)\,\nabla_{\boldsymbol{\theta}_{C}}\mathcal{L}_{k}\Big].$
This correction limits the influence of outdated PQC updates while retaining the throughput benefits of asynchronous aggregation for classical components. The triplet $(\alpha,\eta,w_{k})$, together with optional natural gradient preconditioning, forms the core control space for adapting the hybrid federated optimization loop to varying noise models, hardware availability, and task difficulty.
@@ 6.3 Quantum-Inspired Classical Algorithms
With the introduction of quantum neural computation by [QNC], QNN received significant attention from the QFL community. This pioneering theoretical framework is a crucial advancement in the integration of quantum computation concepts with traditional neural network models, setting the stage for subsequent studies of quantum-inspired classical algorithms. Building on this foundation, numerous quantum-enhanced classical techniques have been developed, each targeting particular computational constraints of classical systems while utilizing quantum benefits.
Building on the challenges of quantum circuit stability, FedQNN [QFL2] presents a more generalized federated quantum framework for various datasets, including genomics and healthcare. A critical strength of the framework is its adaptability to different quantum hardware architectures, evidenced by evaluations on IBM QPUs. The authors in [QFL-QCNN] introduced a novel FL framework using QCNNs for privacy-preserving collaborative learning, particularly suited for non-IID medical data distributions. The novelty lies in the use of quantum circuits to enhance feature extraction and classification within a federated setting. The development of the Quantum Long Short-Term Memory (QLSTM) [QFL-QLSTM] network represents a significant advancement aimed at overcoming the challenges faced by classical LSTMs [sadhwani20245g], especially in managing long-term dependencies in time-series data. Further, Quantum Fuzzy Neural Networks (QFuzzyNNs) [QFL-QFuzzyNN], inspired by tensor network methodologies, present a unique combination of quantum computing principles and fuzzy logic. By utilizing quantum superposition and entanglement, QFuzzyNNs conduct fuzzy inference, enhancing decision-making capabilities even with incomplete information [8651334]. In Fig. 9, we present the advantages and disadvantages of various neural networks when combined with quantum capabilities.
@@
On a slightly different track, QAOA [QAOA] has simultaneously established itself as a potent method to address combinatorial optimization problems. QAOA functions by alternating between quantum and classical optimization stages, enabling it to effectively search through vast solution spaces typically beyond the reach of classical algorithms. The principal innovation of QAOA is its application of quantum circuits for state evolution, followed by classical updates to refine the variational parameters [QFL-QAOA]. This hybrid method has demonstrated remarkable potential in solving the MaxCut problem, consistently surpassing classical algorithms in both convergence speed and solution quality. The capacity of QAOA to provide near-optimal solutions for NP-hard problems underscores its importance in diverse optimization tasks, ranging from logistics [azzaoui2021quantum, dalal2024digitized] to cryptography [phab2022first, kashapogu2024exploring]. Recently, researchers have explored the application of quantum tensors for cyber-physical-social systems [QFL-Tensor, QFL-Tensor2] and revealed that tensor decomposition techniques can substantially decrease communication overhead in distributed learning settings.
An inspection of $\mathrm{QUE}$ for quantum-inspired classical algorithms reveals varied occupancy levels tied to circuit depth and qubit usage. FedQNN’s generalized framework ($\mathrm{Q}_{alloc}=5$, $L=3$) attains $\mathrm{QUE}\approx 1.0$, making it well suited to asynchronous updates that exploit its high utilization and low-latency gradient pushes [QFL2]. The QCNN-based privacy model ($\mathrm{Q}_{alloc}=4$, $L=2$) yields $\mathrm{QUE}=1.0$ and favors synchronous aggregation to ensure consistent feature extraction across non-IID medical nodes [QFL-QCNN]. QLSTM’s deeper memory-enhancing circuits ($\mathrm{Q}_{alloc}=4$, $L=4$) report $\mathrm{QUE}\approx 1.0$ and employ specialized asynchronous updates to balance sequence processing with qubit coherence constraints [QFL-QLSTM]. Despite these strengths, tensor-network hybrids ($\mathrm{Q}_{alloc}=4$, $L=5$) show slightly reduced QUE ($\approx 0.80$) and require hybrid classical–quantum fusion to reconcile their expressive power with aggregation consistency [QFL-Tensor2]. Limits unique to these hybrid frameworks include the overhead of repeated state re-preparation for classical data encoding.